OpenStudy (anonymous):

I need to find a polynomial for the perimeter and the area y+5 on the top and y on the side

OpenStudy (anonymous):

Perimeter P=2w+2l, area A=w*l w=y, l=y+5 P=2y+2(y+5)=2y+2y+10=4y+10 A=y*(y+5)=y^2+5y

OpenStudy (anonymous):

so what is my answer?

OpenStudy (anonymous):

 Area = top x side perimeter = top + bottom + left + right